AutoZone sells a regular car battery for $700 and heavy-duty model for $1100. If it sold 40 batteries yesterday for a total of $34000. Indicate how many of each of the type of batteries and indicate the dollar value sales of each type.
Answer by Theo(13342)

x + y = 40
700x + 1100y = 34000
multiply both sides of the first equation by 700 and leave the second equation alone to get:
700x + 700y = 28000
700x + 1100y = 34000
subtract the first equation from the second to get:
400y = 6000
solve for y to get:
y = 15
that makes x = 25
700 * 25 + 1100 * 15 = 3400, confirming the values of x and y are correct.
your solution is that 25 batteries at 700 each and 15 batteries at 1100 each were sold.
total sales were 25 * 700 = 17500 for the regular batters + 15 * 1100 = 16500 for the heavy duty batteries = total sales of 34000